OpenStudy (anonymous):

I have photos from a multiple choice test that I NEED to pass in order to get into college

OpenStudy (anonymous):

@mathstudent55 help?

OpenStudy (anonymous):

I have photos from a test that I NEED to pass in order to get into college!!

OpenStudy (anonymous):

Its multiple choice too

OpenStudy (anonymous):

@rosedewittbukater are you helping me?! THANK UUUUU

OpenStudy (rosedewittbukater):

For the first one you could use the pythagorean theorem. a^2 + b^2 = c^2 and see if the 2 smaller sides squared added together equals the largest side squared. For the second one get both equations into slope intercept form y=mx+b and see if the m (the slope) of both equations are the same. The third one, get the 2 back inside the radical by ^2 ing it and multiplying by 7. Then square all of the numbers (the square root of 28 squared is 28) and add together the 2 smaller numbers to see if they equal the bigger number. An octagonal prism has 24 edges. For the last one on the first picture I think its true because it can have some triangular faces.

OpenStudy (rosedewittbukater):

I'm only a freshman but I take honors algebra 2 so I'll see how much I can help. So far it doesn't seem too hard.

OpenStudy (anonymous):

I got true, true, false, true, false

OpenStudy (rosedewittbukater):

For the first page?

OpenStudy (anonymous):

yes, I know that #4 is false now

OpenStudy (anonymous):

I double checked

OpenStudy (rosedewittbukater):

27^2 + 31^2 = 48^2 729 + 961 = 2304 1690 = 2304 That's what I got for number 1, so I think its false.

OpenStudy (anonymous):

oh okay

OpenStudy (rosedewittbukater):

number 2 2x+5y = 20 5y = -2x + 20 y = -2/5x + 10 2x - 5y = 35 -5y = -2x + 35 y = 2/5x + 7 Since 2/5 is not the same as -2/5, its not parallel.

OpenStudy (rosedewittbukater):

So 2 is false also.

OpenStudy (anonymous):

Hey, I really still need a lot of help, but I have to go babysit soon, Ill be back on around 4 if you could still help? if you have gmail maybe we can chat?

OpenStudy (rosedewittbukater):

I might be gone by then but you can message me on here if that works.

OpenStudy (anonymous):

the whole test is multiple choice too

OpenStudy (rosedewittbukater):

Oh ok good

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

Ill probably just ask a question and post all of the pictures i have in it and tag you

OpenStudy (abb0t):

If you're caught cheating on a test in college, you won't even be in college at all.

OpenStudy (anonymous):

dude, its one semester of geometry i failed that im making up

OpenStudy (rosedewittbukater):

ok that's fine. I'll try to explain it so you understand it better because algebra 2 is way worse and more confusing especially if you don't get the basics. Just warning you.

OpenStudy (anonymous):

I already passed alg 2!

OpenStudy (anonymous):

This is just a final test that I need to pass so I can make up a grade

OpenStudy (anonymous):

I need to get all the answers by tonight because the test is tomorrow

OpenStudy (anonymous):

I have to go! bye!

OpenStudy (rosedewittbukater):

Oh ok cool! Good luck! I still have homework so I won't message you right away but I'll try to help as much as I can. Is this the actual test?